Show that the orbit nearest to the nucleus in the hydrogen atom accommodates one electron waves. Given that the energy of the electron in the orbit is 13.6 eV, h=6.625x10-34Js, m=9.11x10-31kg, Bohr radius r1 =5.292x10-11 m.


1
Expert's answer
2021-03-04T11:56:07-0500

The angular momentum for the electron in the first Bohr orbit is


"mvr_0=\frac{h}{2\pi}."

Rearrange and notice there is de Broglie wave:


"2\pi r_0=\frac h{mv}=\lambda."

Substitute the values and get


"\lambda=2\pi r_0=333\text{ pm}."

Thus, the orbit nearest to the nucleus in the hydrogen atom accommodates one electron wave.